OpenStudy (anonymous):

A box with gift inside was totally wrapped in beautiful wrapper. if the box is 20cm by 15cm by 12cm, how wide is the wrapper used excluding the overlap?

Directrix (directrix):

The surface area of the rectangular solid is the perimeter of the base time the height of the box plus twice the area of the two bases. SA = 2*(20 + 15) * 12 + 2*(20*15) SA = 2*(35)*12 + 2*(20*15) = 1440 square cm. ----- Once you learn to use the formula, there is an online surface area calculator for rectangular solids at the following link: http://www.ajdesigner.com/phpgeometricformulas/rectangular_solid_area_a.php

OpenStudy (anonymous):

it's okay thank you